OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now. If x represents Sarah's age now, which of the following expressions represents Ralph's age in six years? 2x 6x 2x + 6 3x + 6

OpenStudy (anonymous):

Let r = ralphs age, and x=saras age in six years, ralph will be r+6 years old, and sara will be x+6 years old. but since ralph will be two times as old as sara, (r+6)=2(x+6) you just have to simplify that equation and solve for r. i hope that helped!

OpenStudy (jdoe0001):

ralph = r sara = s ralph is TODAY, 3 times older, so if sara is "s" old, ralph is 3s r = 3s in 6 years, ralph will be 2 times as old as sara THEN, in 6 years ralph will be TODAY's age plus 6, so r + 6 so if sara is "s" old, r+ 6 = 2s r + 6 = 2s

OpenStudy (anonymous):

i agree completely with @jdoe0001